OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x + 3y = 4 x - 2y = -5 @eSpeX

OpenStudy (anonymous):

@cshalvey @robtobey

OpenStudy (espex):

You can use elimination to quickly solve for 'x', however using substitution will work as well.

OpenStudy (anonymous):

How would I do the elimination way?

OpenStudy (espex):

Elimination involved "eliminating" one of the variables, in this case, y. So you multiply each equation by a number that will allow you to get the 'y' values equal.

OpenStudy (primeralph):

multiply the second by -2 and add both equations

OpenStudy (espex):

What do you get if you multiply the two 'y' coefficients together?

OpenStudy (anonymous):

6y, right?

OpenStudy (espex):

Right, so what would you need to multiply into equation 1 to have 6y and in equation 2 to get -6y?

OpenStudy (anonymous):

I have no clue, this is confusing.

OpenStudy (espex):

In equation 1, you have 3y, so you need to multiply by 2 to get 6y. You must multiply the entire equation to keep from changing it, so (2)(2x+3y=4) => (2*2x)+(2*3y)=(2*4) to give you 4x+6y=8 You do the other one.

OpenStudy (anonymous):

Do the same with multiplying by 2 ?

OpenStudy (espex):

If 2 times the 'y' will give you a 6

OpenStudy (anonymous):

Oh, ok. Maybe I should use 3 then.

OpenStudy (espex):

Exactly, good job.

OpenStudy (anonymous):

3x - 6y = -15

OpenStudy (espex):

Excellent. Now you want to add these two equations together, so x+x, y+y, and total + total: 4x +6y = 8 3x -6y = -15 And your resulting equation will allow you to solve for 'x' and get the answer to this problem. :)

OpenStudy (anonymous):

7x + 12y = -7 ?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

6y + (-6y) is not 12y

OpenStudy (anonymous):

Oh, I didn't notice it was -6y

OpenStudy (anonymous):

Is the answer -1

jimthompson5910 (jim_thompson5910):

x = -1, yep

jimthompson5910 (jim_thompson5910):

use that to find y

OpenStudy (anonymous):

All I needed was x.

OpenStudy (anonymous):

But thanks.

jimthompson5910 (jim_thompson5910):

oh true, but it doesn't hurt to find y (for practice)

OpenStudy (anonymous):

True. I'll write it down and try it when I finish these next questions.